# Partial Identification Using Neural Causal Models with Continuous Treatment

### v2

@@ Abstract
Recent progress in Neural Causal Models (NCMs) showcased how identification and partial identification of causal effects can be automatically carried out via training of neural generative models that respect the constraints encoded in a given causal graph [52, 3]. However, formal consistency of these methods has only been proven for the case of discrete variables or only for linear causal models. In this work, we prove the consistency of partial identification via NCMs in a general setting with both continuous and categorical variables. Further, our results highlight the impact of the design of the underlying neural network architecture in terms of depth and connectivity as well as the importance of applying Lipschitz regularization in the training phase. In particular, we provide a counterexample showing that without Lipschitz regularization this method may not be asymptotically consistent. Our results are enabled by new results on the approximability of Structural Causal Models (SCMs) via neural generative models, together with an analysis of the sample complexity of the resulting architectures and how that translates into an error in the constrained optimization problem that defines the partial identification bounds.
@@ 1 Introduction
Identifying causal quantities from observational data is an important problem in causal inference which has wide applications in economics [1], social science [20], health care [34, 18], and recommendation systems [9]. One common approach is to transform causal quantities into statistical quantities using the ID algorithm [49] and deploy general purpose methods to estimate the statistical quantity. However, in the presence of unobserved confounding, typically the causal quantity of interest will not be point-identified by observational data, unless special mechanisms are present in the data generating process (e.g. instruments, unconfounded mediators, proxy controls). In the presence of unobserved confounding, the ID algorithm will typically fail to return a statistical quantity and declare the causal quantity as non-identifiable.
One remedy to this problem, which we focus on in this paper, is partial identification, which aims to give informative bounds for causal quantities based on the available data. At a high level, partial identification bounds can be defined as follows: find the maximum and the minimum value that a target causal quantity can take, among all Structural Causal Models (SCMs) that give rise to the same observed data distribution and respect the given causal graph (as well as any other structural constraints that one is willing to impose). Note that in the presence of unobserved confounding, there will typically exist many structural mechanisms that could give rise to the same observational distribution but have vastly different counterfactual distributions. Hence, partial identification can be formulated as solving a max and a min optimization problem [3]
@@
where $\theta(\mathcal{M})$ is the causal quantity of interest, $\mathcal{M}^{*}$ is the true model, $\bm{V}$ is the set of observed nodes, $P^{\mathcal{M}}(\bm{V})$ is the distribution of $\bm{V}$ in SCM $\mathcal{M}$, $\mathcal{C}$ is a collection of causal models and $\mathcal{G}_{\mathcal{M}}$ is the causal graph of $\mathcal{M}$ (see Section 2 for formal definitions). Two recent lines of work explore the optimization approach to partial identification. The first line deals with discrete Structure Causal Models (SCMs), where all observed variables are finitely supported. In this case, (1.1) becomes a Linear Programming (LP) or polynomial programming problem and tight bounds can be obtained [37, 4, 5, 27, 44, 45, 40, 8, 54, 56, 14]. The second line of work focuses on continuous models and explores ways of solving (1.1) in continuous settings using various techniques [24, 32, 29, 38, 3, 41].
@@
Recently, Xia et al. [51] formalized the connection between SCMs and generative models (see also [33] for an earlier version of a special case of this connection). This work showcased that SCMs can be interpreted as neural generative models, namely Neural Causal Models (NCMs), that follow a particular architecture that respects the constraints encoded in a given causal graph. Hence, counterfactual quantities of SCMs can be learned by optimizing over the parameters of the underlying generative models. However, there could be multiple models that lead to the same observed data distribution, albeit have different counterfactual distributions. Xia et al. [51] first analyze the approximation power of NCMs for discrete SCMs and employ the max/min approach to verify the identifiability of causal quantities, without the need to employ the ID algorithm. Balazadeh et al. [3] and Hui et al. [30], extend the method in [51] and re-purpose it to perform partial identification by solving the min and max problem in the partial identification formulation over neural causal models.
@@
However, for SCMs with general random variables and functional relationships, the approximation error and consistency of this optimization-based approach to partial identification via NCMs has not been established. In particular, two problems remain open. First, given an arbitrary SCM, it is not yet known if we can find an NCM that produces approximately the same intervention distribution as the original one. Although Xia et al.[51] show it is possible to represent any discrete SCM by an NCM, their construction highly relies on the discrete assumption and cannot be directly generalized to the general case. Moreover, Xia et al. [51] use step functions as the activation function in their construction, which may create difficulties in the training process since step functions are discontinuous with a zero gradient almost everywhere.
@@
Second, since we only have access to $n$ samples from the true distribution, $P^{\mathcal{M}}(\bm{V})$, we need to replace the constraints in (1.1) with their empirical version that uses the empirical distribution of samples $P_{n}^{\mathcal{M^{*}}}(\bm{V})$ in place of the population distribution and looks for NCMs, whose implied distribution lies within a small distance from the empirical distribution. Moreover, even the NCM distribution is typically only accessible through sampling, hence we will need to generate $m_{n}$ samples from the NCM and use the empirical distribution of the $m_{n}$ samples from the NCM in place of the true distribution implied by the NCM. Thus, in practice, we will use a constraint of the form $d(P_{n}^{\mathcal{M^{*}}}(\bm{V}),P_{m_{n}}^{\mathcal{M}}(\bm{V}))\leqslant\alpha_{n}$, where $d$ is some notion of distribution distance and $\alpha_{n}$ accounts for the sampling error. It is not clear that this approach is consistent, converges to the correct partial identification bounds, when the sample size $n$ grows to infinity. Balazadeh et al. [3] only show the consistency of this approach for linear SCMs. Consistency results concerning more general SCMs is still lacking in the neural causal literature.
@@
  • We show that under suitable regularity assumptions, given any Lipschitz SCM, we can approximate it using an NCM such that the Wasserstein distance between any interventional distribution of the NCM and the original SCM is small. Each random variable of the SCM is allowed to be continuous or categorical. We specify two architectures of the Neural Networks (NNs) that can be trained using common gradient-based optimization algorithms (Theorem 3.5, Theorem 3.8 and Corollary 3.9).
  • To construct the NCM approximation, we develop a novel representation theorem of probability measures (Proposition 3.7) that may be of independent interest. Proposition 3.7 implies that under certain assumptions, probability distributions supported on the unit cube can be simulated by pushing forward a multivariate uniform distribution.
@@
There exists a rich literature on partial identification of average treatment effects (ATE) [37, 4, 5, 27, 44, 45, 40, 8, 54, 56, 14, 7, 39, 26]. Balke and Pearl [4, 5] first give an algorithm to calculate bounds on the ATE in the Instrumental Variable (IV) setting. They show that regardless of the exact distribution of the latent variables, it is sufficient to consider discrete latent variables as long as all endogenous variables are finitely supported. Moreover, they discover that (1.1) is an LP problem with a closed-form solution. This LP-based technique was generalized to several special classes of SCMs [44, 8, 27, 46]. For general discrete SCMs, [14, 56, 55] consider transforming the problem (1.1) into a polynomial programming problem. Xia et al. [51] discover the connection between generative models and SCMs. They show that NCMs are expressive enough to approximate discrete SCMs. By setting $\mathcal{C}$ in problem (1.1) to be the collection of NCMs, they apply NCMs for identification and estimation.
For causal models with continuous random variables, the constraints in (1.1) become integral equations, which makes the problem more difficult. One approach is to discretize the constraints. [24] uses stochastic process representation of the causal model in the continuous IV setting and transforms the problem into a semi-infinite program. [32] relaxes the constraints to finite moment equations and solves the problem by the Augmented Lagrangian Method (ALM). The other approach is to use generative models to approximate the observational distribution and use some metric to measure the distance between distributions. [33] first propose to use GAN to generate images. Later, [29] uses Wasserstein distance in the constraint and transforms the optimization problem into a min and max problem. Similarly, [3] solves the optimization problem using Sinkhorn distance to avoid instability during training. They propose to estimate the Average Treatment Derivative (ATD) and use ATD to obtain a bound on the ATE. They also prove the consistency of the proposed estimator for linear SCMs. [41] uses a linear combination of basis functions to approximate response functions. [26] uses sieve to solve the resulting optimization problem in the IV setting. [19] proposes a neural network framework for sensitivity analysis under unobserved confounding.
@@
In Section 2, we introduce the notations and some basic concepts used throughout the paper. Next, in Section 3, we demonstrate how to construct an NCM so that they can approximate a given SCM arbitrarily well. Two kinds of architecture are given along with an approximation error analysis. In Section 4, we highlight the importance of Lipschitz regularization by giving a counterexample that is not consistent. Then, leveraging the previous approximation results, we are able to prove the consistency of this approach under regularization. Finally, we compare our method with the traditional polynomial programming method empirically in Section 4.1.
@@ 2 Preliminary
When we write $\text{Pa}(V_{i})$, we refer to the parents of $V_{i}$ in the $\mathcal{G}_{0}$. To connect with the literature, the causal graph we define is a kind of Acyclic Directed Mixed Graph (ADMG), which is often used to represent SCMs with unobserved variables [49]. Note that we allow one latent variable to enter several nodes, which differs from the common definition. We use $n_{U}$ and $n_{V}$ to denote the number of latent variables and observable variables. Let $\bm{T}\subset\bm{V}$ be a set of treatment variables. The goal is to estimate causal quantities under a given intervention $\bm{T}=\bm{t}$. Formally, the structural equations of the intervened model are
@@
We denote $V_{i}(\bm{t})$ to be the value of $V_{i}$ under the intervention $\bm{T}=\bm{t}$ and $P^{\mathcal{M}}(\bm{V}(\bm{t}))$ to be the distribution of $\bm{V}(\bm{t})$ in $\mathcal{M}$. The notion of a $C^{2}$ component [51] is defined as follows.
@@
The definition of NCMs we use is slightly different from that in [52] because we need to deal with mixed variables in our models. In (1.1), we usually take $\mathcal{C}$ to be the set of all SCMs. However, it is difficult to search over all SCMs since (1.1) becomes an infinite-dimensional polynomial programming problem. As an alternative, we can search over all NCMs. One quantity of common interest in causal inference is the Average Treatment Effect (ATE).
@@
Partial identification can be formulated as estimating the solution to the optimization problems (1.1) [3]. The max and min values $\overline{F}$ and $\underline{F}$ define the interval $[\underline{F},\overline{F}]$ which is the smallest interval we can derive from the observed data without additional assumptions. In particular, if $\overline{F}=\underline{F}$, then the causal quantity is point-identified.
Notations. We use $\|\cdotp\|,\|\cdot\|_{\infty}$ for the $1$-norm and $\infty$-norm and $[n]$ for $\{1,\cdots,n\}$. Bold letters represent sets of random variables. We let $\mathcal{F}_{L}(K_{1},K_{2})$ be the class of Lipschitz $L$-continuous functions $f:K_{1}\rightarrow K_{2}$. We may omit the domain and use $\mathcal{F}_{L}$ when the domain is clear from context. Let $\mathcal{H}(K_{1},K_{2})$ be the set of homeomorphisms from $K_{1}$ to $K_{2}$, i.e., injective and continuous maps in both directions. We define $\epsilon(\mathcal{F}_{1},\mathcal{F}_{2})=\sup_{f_{2}\in\mathcal{F}_{2}}\inf_{f_{1}\in\mathcal{F}_{1}}\|f_{2}-f_{1}\|$ for function classes $\mathcal{F}_{1},\mathcal{F}_{2}$. We use standard asymptotic notation $O(\cdot),\Omega(\cdot)$. Given a measure $\mu$ on $\mathbb{R}^{d_{1}}$ and a measurable function $f:\mathbb{R}^{d_{1}}\rightarrow\mathbb{R}^{d_{2}}$, the push-forward measure $f_{\#}\mu$ is defined as $f_{\#}\mu(B)=\mu\left(f^{-1}(B)\right)$ for all measurable sets $B$. We use $P(X)$ to represent the distribution of random variable $X$. Let $\Delta_{n}=\{(p_{1},\cdots,p_{n}):\sum_{i=1}^{n}p_{i}=1,p_{i}\geqslant 0\}$ be the probability simplex. We use $\text{Categorical}(\bm{p}),\bm{p}\in\Delta_{n}$ to represent categorical distribution with event probability $\bm{p}$. We let $W(\cdot,\cdot)$ be the Wasserstein-1 distance and $S_{\lambda}(\mu,\nu)$ be the Sinkhorn distance [12] with regularization parameter $\lambda>0$.

### v3

# Consistency of Neural Causal Partial Identification

@@ Abstract
Recent progress in Neural Causal Models (NCMs) showcased how identification and partial identification of causal effects can be automatically carried out via training of neural generative models that respect the constraints encoded in a given causal graph xia2022neural ; balazadehPartialIdentificationTreatment2022 . However, formal consistency of these methods has only been proven for the case of discrete variables or only for linear causal models. In this work, we prove the consistency of partial identification via NCMs in a general setting with both continuous and categorical variables. Further, our results highlight the impact of the design of the underlying neural network architecture in terms of depth and connectivity as well as the importance of applying Lipschitz regularization in the training phase. In particular, we provide a counterexample showing that without Lipschitz regularization this method may not be asymptotically consistent. Our results are enabled by new results on the approximability of Structural Causal Models (SCMs) via neural generative models, together with an analysis of the sample complexity of the resulting architectures and how that translates into an error in the constrained optimization problem that defines the partial identification bounds.
@@ 1 Introduction
Identifying causal quantities from observational data is an important problem in causal inference which has wide applications in economics angrist1995identification , social science freedman2010statistical , health care lotfollahi2021compositional ; frauen2023estimating , and recommendation systems bottou2013counterfactual . One common approach is to transform causal quantities into statistical quantities using the ID algorithm tianTestableImplicationsCausal2002 and deploy general purpose methods to estimate the statistical quantity. However, in the presence of unobserved confounding, typically the causal quantity of interest will not be point-identified by observational data, unless special mechanisms are present in the data generating process (e.g. instruments, unconfounded mediators, proxy controls). In the presence of unobserved confounding, the ID algorithm will typically fail to return a statistical quantity and declare the causal quantity as non-identifiable.
One remedy to this problem, which we focus on in this paper, is partial identification, which aims to give informative bounds for causal quantities based on the available data. At a high level, partial identification bounds can be defined as follows: find the maximum and the minimum value that a target causal quantity can take, among all Structural Causal Models (SCMs) that give rise to the same observed data distribution and respect the given causal graph (as well as any other structural constraints that one is willing to impose). Note that in the presence of unobserved confounding, there will typically exist many structural mechanisms that could give rise to the same observational distribution but have vastly different counterfactual distributions. Hence, partial identification can be formulated as solving a max and a min optimization problem balazadehPartialIdentificationTreatment2022
@@
where $\theta(\mathcal{M})$ is the causal quantity of interest, $\mathcal{M}^{*}$ is the true model, $\bm{V}$ is the set of observed nodes, $P^{\mathcal{M}}(\bm{V})$ is the distribution of $\bm{V}$ in SCM $\mathcal{M}$, $\mathcal{C}$ is a collection of causal models and $\mathcal{G}_{\mathcal{M}}$ is the causal graph of $\mathcal{M}$ (see Section 2 for formal definitions). Two recent lines of work explore the optimization approach to partial identification. The first line deals with discrete Structure Causal Models (SCMs), where all observed variables are finitely supported. In this case, (1.1) becomes a Linear Programming (LP) or polynomial programming problem and tight bounds can be obtained manski1990nonparametric ; balkeCounterfactualProbabilitiesComputational1994 ; balkeBoundsTreatmentEffects1997 ; heckmanInstrumentalVariablesSelection2001 ; ramsahaiCausalBoundsObservable2012 ; richardsonNonparametricBoundsSensitivity2014 ; milesPartialIdentificationPure2015 ; bonetInstrumentalityTestsRevisited2013 ; zhangBoundingCausalEffects2021 ; zhangPartialCounterfactualIdentification2021a ; duarteAutomatedApproachCausal2021 . The second line of work focuses on continuous models and explores ways of solving (1.1) in continuous settings using various techniques gunsiliusPathsamplingMethodPartially2020 ; kilbertusClassAlgorithmsGeneral2020 ; huGenerativeAdversarialFramework2021 ; maoGenerativeInterventionsCausal2021 ; balazadehPartialIdentificationTreatment2022 ; padhStochasticCausalProgramming2023 .
@@
Recently, Xia et al. xiaCausalNeuralConnectionExpressiveness2022 formalized the connection between SCMs and generative models (see also kocaoglu2017causalgan for an earlier version of a special case of this connection). This work showcased that SCMs can be interpreted as neural generative models, namely Neural Causal Models (NCMs), that follow a particular architecture that respects the constraints encoded in a given causal graph. Hence, counterfactual quantities of SCMs can be learned by optimizing over the parameters of the underlying generative models. However, there could be multiple models that lead to the same observed data distribution, albeit have different counterfactual distributions. Xia et al. xiaCausalNeuralConnectionExpressiveness2022 first analyze the approximation power of NCMs for discrete SCMs and employ the max/min approach to verify the identifiability of causal quantities, without the need to employ the ID algorithm. Balazadeh et al. balazadehPartialIdentificationTreatment2022 and Hui et al. hu2021generative , extend the method in xiaCausalNeuralConnectionExpressiveness2022 and re-purpose it to perform partial identification by solving the min and max problem in the partial identification formulation over neural causal models.
@@
However, for SCMs with general random variables and functional relationships, the approximation error and consistency of this optimization-based approach to partial identification via NCMs has not been established. In particular, two problems remain open. First, given an arbitrary SCM, it is not yet known if we can find an NCM that produces approximately the same intervention distribution as the original one. Although Xia et al.xiaCausalNeuralConnectionExpressiveness2022 show it is possible to represent any discrete SCM by an NCM, their construction highly relies on the discrete assumption and cannot be directly generalized to the general case. Moreover, Xia et al. xiaCausalNeuralConnectionExpressiveness2022 use step functions as the activation function in their construction, which may create difficulties in the training process since step functions are discontinuous with a zero gradient almost everywhere.
@@
Second, since we only have access to $n$ samples from the true distribution, $P^{\mathcal{M}}(\bm{V})$, we need to replace the constraints in (1.1) with their empirical version that uses the empirical distribution of samples $P_{n}^{\mathcal{M^{*}}}(\bm{V})$ in place of the population distribution and looks for NCMs, whose implied distribution lies within a small distance from the empirical distribution. Moreover, even the NCM distribution is typically only accessible through sampling, hence we will need to generate $m_{n}$ samples from the NCM and use the empirical distribution of the $m_{n}$ samples from the NCM in place of the true distribution implied by the NCM. Thus, in practice, we will use a constraint of the form $d(P_{n}^{\mathcal{M^{*}}}(\bm{V}),P_{m_{n}}^{\mathcal{M}}(\bm{V}))\leqslant\alpha_{n}$, where $d$ is some notion of distribution distance and $\alpha_{n}$ accounts for the sampling error. It is not clear that this approach is consistent, converges to the correct partial identification bounds, when the sample size $n$ grows to infinity. Balazadeh et al. balazadehPartialIdentificationTreatment2022 only show the consistency of this approach for linear SCMs. Consistency results concerning more general SCMs is still lacking in the neural causal literature.
@@
  • We show that under suitable regularity assumptions, given any Lipschitz SCM, we can approximate it using an NCM such that the Wasserstein distance between any interventional distribution of the NCM and the original SCM is small. Each random variable of the SCM is allowed to be continuous or categorical. We specify two architectures of the Neural Networks (NNs) that can be trained using common gradient-based optimization algorithms (Theorem 3.6 and Corollary 3.7).
  • To construct the NCM approximation, we develop a novel representation theorem of probability measures (Proposition B.2) that may be of independent interest. Proposition B.2 implies that under certain assumptions, probability distributions supported on the unit cube can be simulated by pushing forward a multivariate uniform distribution.
@@
There exists a rich literature on partial identification of average treatment effects (ATE) manski1990nonparametric ; balkeCounterfactualProbabilitiesComputational1994 ; balkeBoundsTreatmentEffects1997 ; heckmanInstrumentalVariablesSelection2001 ; ramsahaiCausalBoundsObservable2012 ; richardsonNonparametricBoundsSensitivity2014 ; milesPartialIdentificationPure2015 ; bonetInstrumentalityTestsRevisited2013 ; zhangBoundingCausalEffects2021 ; zhangPartialCounterfactualIdentification2021a ; duarteAutomatedApproachCausal2021 ; bellot2023towards ; melnychukPartialCounterfactualIdentification2023a ; han2024computational . Balke and Pearl balkeCounterfactualProbabilitiesComputational1994 ; balkeBoundsTreatmentEffects1997 first give an algorithm to calculate bounds on the ATE in the Instrumental Variable (IV) setting. They show that regardless of the exact distribution of the latent variables, it is sufficient to consider discrete latent variables as long as all endogenous variables are finitely supported. Moreover, they discover that (1.1) is an LP problem with a closed-form solution. This LP-based technique was generalized to several special classes of SCMs ramsahaiCausalBoundsObservable2012 ; bonetInstrumentalityTestsRevisited2013 ; heckmanInstrumentalVariablesSelection2001 ; sachsGeneralMethodDeriving2021 . For general discrete SCMs, duarteAutomatedApproachCausal2021 ; zhangPartialCounterfactualIdentification2021a ; zhangNonparametricMethodsPartial2021 consider transforming the problem (1.1) into a polynomial programming problem. Xia et al. xiaCausalNeuralConnectionExpressiveness2022 discover the connection between generative models and SCMs. They show that NCMs are expressive enough to approximate discrete SCMs. By setting $\mathcal{C}$ in problem (1.1) to be the collection of NCMs, they apply NCMs for identification and estimation.
For causal models with continuous random variables, the constraints in (1.1) become integral equations, which makes the problem more difficult. One approach is to discretize the constraints. gunsiliusPathsamplingMethodPartially2020 uses stochastic process representation of the causal model in the continuous IV setting and transforms the problem into a semi-infinite program. kilbertusClassAlgorithmsGeneral2020 relaxes the constraints to finite moment equations and solves the problem by the Augmented Lagrangian Method (ALM). The other approach is to use generative models to approximate the observational distribution and use some metric to measure the distance between distributions. kocaoglu2017causalgan first propose to use GAN to generate images. Later, huGenerativeAdversarialFramework2021 uses Wasserstein distance in the constraint and transforms the optimization problem into a min and max problem. Similarly, balazadehPartialIdentificationTreatment2022 solves the optimization problem using Sinkhorn distance to avoid instability during training. They propose to estimate the Average Treatment Derivative (ATD) and use ATD to obtain a bound on the ATE. They also prove the consistency of the proposed estimator for linear SCMs. padhStochasticCausalProgramming2023 uses a linear combination of basis functions to approximate response functions. han2024computational uses sieve to solve the resulting optimization problem in the IV setting. frauenNeuralFrameworkGeneralized2024 proposes a neural network framework for sensitivity analysis under unobserved confounding.
@@
In Section 2, we introduce the notations and some basic concepts used throughout the paper. Next, in Section 3, we demonstrate how to construct an NCM so that they can approximate a given SCM arbitrarily well. Two kinds of architecture are given along with an approximation error analysis. In Section 4, we highlight the importance of Lipschitz regularization by giving a counterexample that is not consistent. Then, leveraging the previous approximation results, we are able to prove the consistency of this approach under regularization. Finally, we compare our method with the traditional polynomial programming method empirically in Section 5.
@@ 2 Preliminary
When we write $\text{Pa}(V_{i})$, we refer to the parents of $V_{i}$ in the $\mathcal{G}_{0}$. To connect with the literature, the causal graph we define is a kind of Acyclic Directed Mixed Graph (ADMG), which is often used to represent SCMs with unobserved variables tianTestableImplicationsCausal2002 . Note that we allow one latent variable to enter several nodes, which differs from the common definition. We use $n_{U}$ and $n_{V}$ to denote the number of latent variables and observable variables. Let $\bm{T}\subset\bm{V}$ be a set of treatment variables. The goal is to estimate causal quantities under a given intervention $\bm{T}=\bm{t}$. Formally, the structural equations of the intervened model are
@@
We denote $V_{i}(\bm{t})$ to be the value of $V_{i}$ under the intervention $\bm{T}=\bm{t}$ and $P^{\mathcal{M}}(\bm{V}(\bm{t}))$ to be the distribution of $\bm{V}(\bm{t})$ in $\mathcal{M}$. The notion of a $C^{2}$ component xiaCausalNeuralConnectionExpressiveness2022 is defined as follows.
@@
The definition of NCMs we use is slightly different from that in xia2022neural because we need to deal with mixed variables in our models. In (1.1), we usually take $\mathcal{C}$ to be the set of all SCMs. However, it is difficult to search over all SCMs since (1.1) becomes an infinite-dimensional polynomial programming problem. As an alternative, we can search over all NCMs. One quantity of common interest in causal inference is the Average Treatment Effect (ATE).
@@
Partial identification can be formulated as estimating the solution to the optimization problems (1.1) balazadehPartialIdentificationTreatment2022 . The max and min values $\overline{F}$ and $\underline{F}$ define the interval $[\underline{F},\overline{F}]$ which is the smallest interval we can derive from the observed data without additional assumptions. In particular, if $\overline{F}=\underline{F}$, then the causal quantity is point-identified.
Notations. We use $\|\cdotp\|,\|\cdot\|_{\infty}$ for the $1$-norm and $\infty$-norm and $[n]$ for $\{1,\cdots,n\}$. Bold letters represent sets of random variables. We let $\mathcal{F}_{L}(K_{1},K_{2})$ be the class of Lipschitz $L$-continuous functions $f:K_{1}\rightarrow K_{2}$. We may omit the domain and use $\mathcal{F}_{L}$ when the domain is clear from context. Let $\mathcal{H}(K_{1},K_{2})$ be the set of homeomorphisms from $K_{1}$ to $K_{2}$, i.e., injective and continuous maps in both directions. We define $\epsilon(\mathcal{F}_{1},\mathcal{F}_{2})=\sup_{f_{2}\in\mathcal{F}_{2}}\inf_{f_{1}\in\mathcal{F}_{1}}\|f_{2}-f_{1}\|$ for function classes $\mathcal{F}_{1},\mathcal{F}_{2}$. We use standard asymptotic notation $O(\cdot),\Omega(\cdot)$. Given a measure $\mu$ on $\mathbb{R}^{d_{1}}$ and a measurable function $f:\mathbb{R}^{d_{1}}\rightarrow\mathbb{R}^{d_{2}}$, the push-forward measure $f_{\#}\mu$ is defined as $f_{\#}\mu(B)=\mu\left(f^{-1}(B)\right)$ for all measurable sets $B$. We use $P(X)$ to represent the distribution of random variable $X$. Let $\Delta_{n}=\{(p_{1},\cdots,p_{n}):\sum_{i=1}^{n}p_{i}=1,p_{i}\geqslant 0\}$ be the probability simplex. We use $\text{Categorical}(\bm{p}),\bm{p}\in\Delta_{n}$ to represent categorical distribution with event probability $\bm{p}$. We let $W(\cdot,\cdot)$ be the Wasserstein-1 distance and $S_{\lambda}(\mu,\nu)$ be the Sinkhorn distance chizat2020faster with regularization parameter $\lambda>0$.